OpenStudy (gabylovesyou):

Fiona decides to join a gym for a minimum of one month to train for a triathlon. The gym charges a joining fee of $100 and a monthly fee of $38. If x represents the number of months that Fiona is a member of the gym, the equation below can be used to determine C, her total membership fee for that duration of time. 100 + 38x = C Fiona has allocated a maximum of $328 to spend on her gym membership. Which number line shows the possible number of months that Fiona can be a member of the gym?

OpenStudy (gabylovesyou):

@jim_thompson5910

OpenStudy (anonymous):

328-100=228/38=6

OpenStudy (anonymous):

so C=328 since it is the maximum she has spent

OpenStudy (gabylovesyou):

how would this be drawn on a number line?

OpenStudy (anonymous):

so you plug it in and you get this: 100+38x=328 38x=228 divide both sides by 38 and you get x=6 x=months so she's been a member 6 months max.

OpenStudy (gabylovesyou):



OpenStudy (anonymous):

that looks right since shading it in makes it equal to.. and it can be x is less than or EQUAL to 6. so yup! u got it :)